A lab technician has one solution that is 60% chlorinated and another that is 40% chlorinated. how much of each solution is needed to make a 100 L solution that is 50% chlorine?
Answer by jorel1380(3719)

m+n=100
.6m+.4n=100(.5)=50
6m+4n=500
4m+4n=400
2m=100
m=50
n=50
You need 50 mls. each of 60% and 40% chlorinated solutions..